# Measuring gravitational lensing time delays with quantum information processing

### v1

@@ I.1 Motivation and background
The rapidly developing field of quantum information technology has various promising applications. In particular, it has been widely accepted that certain computational problems can be solved more efficiently by quantum computers Shor (1999); Grover (1996), and certain signals can be measured more precisely using quantum sensing Giovannetti et al. (2006). Some quantum sensing protocols provide improvement by reducing the number of photons needed, for instance, but quantum technology can also provide other advantages for sensing. An example along these lines is the quantum telescope array proposed by Gottesman, Jennewein, and Croke (2012), a long-baseline optical interferometer scheme for astronomical observation. The proposed telescope provides ultra-high angular resolution (micro-arcsec) using error-corrected quantum communication techniques as the key building block. Gottesman et al. (2012) stimulated various follow-up research in applying quantum information processing techniques to develop novel strategies for optical imaging and astronomical observation Khabiboulline et al. (2019a, b); Liu et al. (2024); Brown et al. (2023).
@@
In this paper, we focus on measuring the time delay between optical signals. This is of great significance in astronomy due to applications to observing gravitational lensing Dyson et al. (1920); Einstein (1936) events. Gravitational lensing occurs when light from a distant source is bent by the gravitational field of an intervening massive object, leading to the formation of multiple images. As a result of this deflection, different light paths associated with the lensed images correspond to different geometric lengths and geodesics, thereby introducing relative time delays in their arrival times Refsdal (1964). These time delays provide a direct and powerful means to measure the mass profile of the lens system, including masses of rogue planets Miret-Roig (2023); Koshimoto et al. (2023); Mróz et al. (2017), isolated black holes Sahu et al. (2022); Kaczmarek et al. (2025), and even the spatial distribution of dark matter, an unknown component of the mass budget of the universe whose contribution dominates over that of visible matter Planck Collaboration et al. (2020).
@@
While measurements of lensing time delays are of great scientific value, it is highly challenging to obtain them in practice. Successful measurements have only been done for strong gravitational lensing systems Biggs et al. (1999); Suyu et al. (2017); Chen et al. (2019), where “strong” means that different images of the source are spatially distinguishable. In strong lensing, the time delay is measured by exploiting variability of the source. For instance, a transient astrophysical event, such as a supernova explosion or a quasar flare, appears in each lensed image at different times due to their light path difference. However, in other lensing systems, such as instances of microlensing, the different images cannot be individually resolved. The time delay of a microlensing event is typically much shorter than the source’s variability timescale, unlike in strong lensing, hence transient signals propagating along two different paths overlap in the recombined light curve, hiding the small time delay between them. Furthermore, the source in a microlensing system is often selected to be absent of well-defined, observable transient events and to remain stable over long baselines. Together, these factors make it infeasible to use techniques for strong lensing to extract time delays in microlensing systems.
@@
Fortunately, there is a different theoretical framework that is designed particularly for measuring microlensing time delays. This framework has multiple observational proposals and one actual implementation. Throughout this paper, we assume there are only two images in the lensing system (this assumption is justified in Section II.1) and denote their time delay by $\Delta t$. References Deguchi and Watson (1986); Peterson and Falk (1991) first realized that a fixed gravitational lensing time delay leads to an oscillatory modulation in the spectrum, with adjacent peaks separated by $1/\Delta t$. This observation laid the theoretical foundation for various follow-up works. Later, Refs. Gould (1992); Ulmer and Goodman (1995); Katz et al. (2018) predicted the existence of lensing events induced by extremely lightweight lensing objects and claimed that their (very short) time delays may create observable frequency-domain oscillations when the photon source is a gamma-ray burst. Similarly, Refs. Eichler (2017); Jow et al. (2020); Wucknitz et al. (2021) discuss lensing delay measurement in the radio wavelength, where they use fast radio bursts (FRBs) as their photon source. There is even one reported experimental attempt using the FRB-based proposal Kader et al. (2022) which enables constraining the abundance of primordial black holes.
@@
However, none of the above proposals provide successful measurement outcomes for any microlensing system, for multiple reasons. First, the femtolensing observations using gamma-ray bursts and most proposals beyond the radio wavelength suffer from the severe finite-source effect. If the source is large, photons from different regions of the source have different $\Delta t$ values. When the time delay uncertainty $\delta_{\Delta t}$ is greater than one period of the carrier frequency, the gravitationally lensed light signal no longer contains any information about the time delay. Such an effect is studied in previous works Eichler (2017); Leung et al. (2025); Sugiyama et al. (2020) and will also be shown information theoretically in our work. Second, the number of photons required in the aforementioned proposals is generally large, while microlensing systems have, on average, lower signal-to-noise ratio (SNR) than strong lensing systems because their dynamic nature limits opportunities to do stable long exposure. To achieve a high SNR, the integration times are sometimes required longer than the lensing event itself and forces researchers to use extremely luminous sources in their designated wavelengths, such as FRBs in radio wavelengths. However, the number of FRB events per day is (empirically) limited to a small value, and the number of gravitationally microlensed FRB events is even lower. Finally, when $\Delta t$ is relatively large, say $\sim 1\,\mathrm{ms}$, the distance between peaks of the spectrum is only $1/\Delta t\sim 1\,\mathrm{kHz}$. Observing such a pattern requires prohibitively high frequency resolution for carrier frequencies higher than those of radio waves.
@@ I.2 Our contribution
Our delay-finding approach relies on a key observation that every photon emitted in a spherical wave takes both paths created by gravitational lensing to reach the Earth in quantum superposition. With this, inspired by the advancement of quantum information science and building upon the intuition of frequency-domain interference from Refs. Deguchi and Watson (1986); Peterson and Falk (1991), we propose a concrete sample-efficient delay measurement approach, Algorithm 1, that provably uses as few photons as possible. In particular, letting $T$ be the upper limit of $\Delta t$, $t_{c}$ be the coherence time of the photon without lensing (which corresponds to the width of the wave packet in the time domain and can be controlled by the bandwidth of the filter we use during observation), and assuming $\Delta t\gg t_{c}$, our method consumes only $O(\log(T/t_{c}))$ photons to measure $\Delta t$ with precision $t_{c}$, while traditional proposals require $O(T/t_{c})$ photons. Note that we work in the photon-starved regime, hence we expect to receive at most one photon per wave packet, and the wave function of each photon is the superposition of two wave packets separated by $\Delta t$ due to the microlensing effect. We also provide a rigorous proof that $\Omega(\log(T/t_{c}))$ is the information-theoretic lower bound, hence our method is optimal. One proof is based on modeling the gravitational lensing system as a communication channel and computing its channel capacity. We also exploit a surprising connection between the delay-finding problem and a well-studied problem in quantum computing, the dihedral hidden subgroup problem. We show that the dihedral hidden subgroup problem can be reduced to our problem, giving an alternative optimality proof in terms of both sample complexity and computational complexity.
@@
The key ingredient for the exponential improvement in our scheme is that our algorithm uses quantum information processing technologies (including single-photon spectrometers, and, depending on the specific implementation of our scheme, quantum memory and digital quantum computation) to perform single-photon frequency-basis measurements. This allows us to sample from a certain distribution determined by the value of $\Delta t$. By feeding these samples into a data-processing procedure inspired by the sample-efficient algorithm for the dihedral hidden subgroup problem Ettinger and Høyer (2000), we can estimate $\Delta t$ in the style of maximum-likelihood estimation.
@@
Implementing our approach involves measuring the frequency of every photon, which requires a broadband high-resolution spectrometer with single-photon sensitivity. The difficulty of the implementation strongly depends on $T$, the upper limit of $\Delta t$ (which is determined by the lensing object) of our interest, because the required frequency resolution is $\sim 1/\Delta t$. Single-photon spectrometers based on dual-combs feature up to $\sim 100\,\mathrm{MHz}$ resolution Coddington et al. (2016); Picqué and Hänsch (2020); Xu et al. (2024); Peng et al. (2025) with $\sim 10\,\mathrm{GHz}$ bandwidth. State-of-the-art spectrometers based on a time lens Mazelanik et al. (2020); Joshi et al. (2022); Mazelanik et al. (2020) even achieve $20\,\mathrm{kHz}$ resolution, but their bandwidth is limited to MHz-level. These results in principle enable measurements of $10\,\mathrm{ns}$ (dual-comb) or up-to-$0.1\,\mathrm{ms}$ (time lens) time delay, corresponding to lensing objects as heavy as brown dwarfs (dual-comb) or primordial black holes of multiple solar masses (time lens), as is explained later in Section II.1. However, one bottleneck for such an observation is that the photon sources we consider are thermal sources emitting broadband signals. The tiny bandwidth of existing single-photon spectrometers may require using prohibitively many of them in parallel. Therefore, long-$\Delta t$ measurements may only be achievable through next-generation single-photon spectrometers; however, for shorter $\Delta t$, which are also of significant interest to astronomy, resolving the frequency of single photons is much less challenging and can potentially be realized by combining existing technologies.
@@
We also propose another version of the delay-finding approach (Algorithm 2) which relies on storing and processing the photon wave function in the time domain and uses a different data processing procedure. We propose to first perform non-demolition frequency measurement on the received photon to localize it to a frequency range of $1/t_{c}$-width. Then, using a quantum information discarding process, one can store the photon in a digital quantum computer in an undersampling manner in the time domain. More specifically, we propose to use a quantum memory that can only distinguish $O(T/t_{c})$ modes, which is far fewer than $\Theta(T\omega_{0})$ modes for sampling at the Nyquist rate (where $\omega_{0}$ is the carrier frequency of the photon). With this approach, one can employ the quantum Fourier transform (QFT) to produce an aliased frequency as the output, which is fed into Algorithm 2 to find $\Delta t$. The connection between delay finding and the dihedral hidden subgroup problem is established through this time-domain version. We show that this version can potentially be implemented by a linear optics system. Going further, digital quantum computing in principle enables compressed storage of the photonic modes with binary encoding of arrival time, which gives an exponential reduction in the resources required.
@@
Our photon-efficient method enables the estimation of microlensing time delays in the optical and infrared (IR) bands. The measurable delay range spans from $10^{-10}$ to $10^{-3}$ seconds depending on the capabilities of the frequency-resolving device. This, in principle, supports observation of many more sources than in radio/gamma wavelengths, and the photons can be lensed by the majority of interesting microlensing systems. However, optical/IR waves oscillate extremely fast, giving a rather stringent requirement on the variance of $\Delta t$ between different photons due to the finite-source effect. This means it is much more difficult to have ingredient (ii). Nevertheless, as part of our solution, we give a concrete use case of our measurement scheme in the optical/IR band that satisfies ingredient (ii), and present a comprehensive analysis of its scientific value and feasibility. Specifically, we consider flares of M-class red dwarfs (M dwarfs). M dwarfs are relatively tiny and cold stars, and a flare is an event in which a small region of the dwarf becomes almost as bright as the whole dwarf in certain passbands. Our analysis shows that, for a significant fraction of flares in M dwarfs, the size of the light-emitting area may be small enough such that the uncertainty in $\Delta t$, denoted by $\delta_{\Delta t}$, is less than one period of the carrier frequency $\sim 10^{-14}$s. Moreover, our scheme not only enables the study of the lensing object, but also yields constraints on the actual spatial size of the flares in M dwarfs, which is currently poorly understood. We also perform a comprehensive analysis of the number of photons we can receive in realistic settings to observe microlensed flares on M dwarfs, taking into account the duration, size, and temperature of the flare, as well as astronomical dust extinction and telescope collecting area. Our result shows that near-term ground-based optical telescopes can achieve sufficiently high signal-to-noise ratio in such observations. Moreover, we improve our algorithms such that photons from temporally and spatially separated flares can be analyzed collectively to infer the average lensing time delay, allowing for potential implementation using existing optical telescopes.
@@
Finally, we discuss the application of our methods to a different task: the calibration of time delays in telescope arrays. Light traveling from a source along different paths picks up a relative delay when observed at different sites. Learning these time delays is important to enable interference of photons arriving at different telescopes. In order to learn the time delays, we map the distributed problem to a lensing-like scenario, where a single detector observes the photons. In particular, we show how to use entanglement to transfer the information across the array to a single site, where we can apply our algorithm. The same compression and storage of photonic information in memory as used in prior work on telescope arrays Khabiboulline et al. (2019a, b) is applicable here, such that our proposal is compatible with that scheme. The benefits over classical techniques are a replacement of long delay lines with memories that keep track of timing, allowing for longer baselines and thus larger resolution, and improved sample efficiency, as provided by our algorithm.
@@
The remainder of the paper is structured as follows. We provide preliminary information in Section II: we give a technical introduction to gravitational lensing and the significance of measuring its time delay (LABEL:{sec:lens_intro}), explain the setup of the delay-finding problem with mathematical and physical rigor (Section II.2), and review a traditional time delay measurement approach using a large number of photons (Section II.3). In Section III, we describe our sample-efficient delay-finding algorithm: we introduce the frequency-domain interference framework via a classical electromagnetism derivation (Section III.1), reproduce the same derivation in the quantum picture for photonic wave functions (Section III.2), use this picture to propose our delay-finding algorithm and analyze its sample complexity (Section III.3), explain the consequence of varied $\Delta t$ values (the finite-source effect) and analyze the how our algorithm performs under this effect (Section III.4), and finally analyze the effect of noise photons (Section III.5) and the lensing magnification (Section III.6). In Section IV, we prove the information-theoretic lower bound for the sample complexity, matching the actual sample complexity of our algorithm. In Section V, we propose the undersampling algorithm using the quantum Fourier transform on a digital quantum computer, and discuss its connection to the dihedral hidden subgroup problem (Section V.3), giving another proof of the optimal sample complexity as well as computational complexity. In Section VI, we discuss possible experimental realizations of our algorithms. In Section VII, we present the astronomical observation plan: we carry out a case study for the example setup for microlensed M dwarf flares with an analysis of its feasibility and scientific value (Section VII.1); we then introduce a modified version of our algorithm to combine photons from different flares and present numerical simulation results (Section VII.2). In Section VIII, we perform a thorough analysis for the robustness of our approach to noises due to the medium between the source and the telescope: we investigate the effect of dust extinction (Section VIII.1 and a detailed proof in Appendix A), astronomical scintillation (Section VIII.2), and atmospheric fluctuation (Section VIII.3). In Section IX, we discuss the application of our delay-finding protocol to calibrating quantum telescope arrays. Finally, we summarize our work and discuss open problems in Section X.

### v2

@@ I.1 Motivation and background
The rapidly developing field of quantum information technology has various promising applications. In particular, it has been widely accepted that certain computational problems can be solved more efficiently by quantum computers [1, 2], and certain signals can be measured more precisely using quantum sensing [3]. Some quantum sensing protocols provide improvement by reducing the number of photons needed, for instance, but quantum technology can also provide other advantages for sensing. An example along these lines is the quantum telescope array proposed by Gottesman, Jennewein, and Croke [4], a long-baseline optical interferometer scheme for astronomical observation. The proposed telescope provides ultra-high angular resolution (micro-arcsec) using error-corrected quantum communication techniques as the key building block. Gottesman et al. [4] stimulated various follow-up research in applying quantum information processing techniques to develop novel strategies for optical imaging and astronomical observation [5, 6, 7, 8].
@@
In this paper, we focus on measuring the time delay between optical signals. This is of great significance in astronomy due to applications to observing gravitational lensing [9, 10] events. Gravitational lensing occurs when light from a distant source is bent by the gravitational field of an intervening massive object, leading to the formation of multiple images. As a result of this deflection, different light paths associated with the lensed images correspond to different geometric lengths and geodesics, thereby introducing relative time delays in their arrival times [11]. These time delays provide a direct and powerful means to measure the mass profile of the lens system, including masses of rogue planets [12, 13, 14], isolated black holes [15, 16], and even the spatial distribution of dark matter, an unknown component of the mass budget of the universe whose contribution dominates over that of visible matter [17].
@@
While measurements of lensing time delays are of great scientific value, it is highly challenging to obtain them in practice. Successful measurements have only been done for strong gravitational lensing systems [18, 19, 20], where “strong” means that different images of the source are spatially distinguishable. In strong lensing, the time delay is measured by exploiting variability of the source. For instance, a transient astrophysical event, such as a supernova explosion or a quasar flare, appears in each lensed image at different times due to their light path difference. However, in other lensing systems, such as instances of microlensing, the different images usually cannot be individually resolved  [^1] Different images in a microlensing system are usually not resolved with seeing-limited imaging; however, long-baseline interferometry (e.g., VLTI/GRAVITY) has now resolved microlensed images in select cases (see e.g. Ref. [105]). Whereas long-baseline interferometers can resolve the micro-images and infer lens parameters from visibilities/closure phases, our method infers the time delay $\Delta t$ from single-photon spectra without resolving the images. The time delay of a microlensing event is typically much shorter than the source’s variability timescale, unlike in strong lensing, hence transient signals propagating along two different paths overlap in the recombined light curve, hiding the small time delay between them. Furthermore, the source in a microlensing system is often selected to be absent of well-defined, observable transient events and to remain stable over long baselines. Together, these factors make it infeasible to use techniques for strong lensing to extract time delays in microlensing systems.
@@
Fortunately, there is a different theoretical framework that is designed particularly for measuring microlensing time delays. This framework has multiple observational proposals and one actual implementation. Throughout this paper, we assume there are only two images in the lensing system (this assumption is justified in Section II.1) and denote their time delay by $\Delta t$. References [22, 23] first realized that a fixed gravitational lensing time delay leads to an oscillatory modulation in the spectrum, with adjacent peaks separated by $1/\Delta t$. This observation laid the theoretical foundation for various follow-up works. Later, Refs. [24, 25, 26] predicted the existence of lensing events induced by extremely lightweight lensing objects and claimed that their (very short) time delays may create observable frequency-domain oscillations when the photon source is a gamma-ray burst. Similarly, Refs. [27, 28, 29] discuss lensing delay measurement in the radio wavelength, where they use fast radio bursts (FRBs) as their photon source. There is even one reported experimental attempt using the FRB-based proposal [30] which enables constraining the abundance of primordial black holes.
@@
However, none of the above proposals provide successful measurement outcomes for any microlensing system, for multiple reasons. First, the femtolensing observations using gamma-ray bursts and most proposals beyond the radio wavelength suffer from the severe finite-source effect. If the source is large, photons from different regions of the source have different $\Delta t$ values. When the time delay uncertainty $\delta_{\Delta t}$ is greater than one period of the carrier frequency, the gravitationally lensed light signal no longer contains any information about the time delay. Such an effect is studied in previous works [27, 31, 32] and will also be shown information theoretically in our work. Second, the number of photons required in the aforementioned proposals is generally large, while microlensing systems have, on average, lower signal-to-noise ratio (SNR) than strong lensing systems because their dynamic nature limits opportunities to do stable long exposure. To achieve a high SNR, the integration times are sometimes required longer than the lensing event itself and forces researchers to use extremely luminous sources in their designated wavelengths, such as FRBs in radio wavelengths. However, the number of FRB events per day is (empirically) limited to a small value, and the number of gravitationally microlensed FRB events is even lower. Finally, when $\Delta t$ is relatively large, say $\sim 1\,\mathrm{ms}$, the distance between peaks of the spectrum is only $1/\Delta t\sim 1\,\mathrm{kHz}$. Observing such a pattern requires prohibitively high frequency resolution for carrier frequencies higher than those of radio waves.
@@ I.2 Our contribution
Our delay-finding approach relies on a key observation that every photon emitted in a spherical wave takes both paths created by gravitational lensing to reach the Earth in quantum superposition. With this, inspired by the advancement of quantum information science and building upon the intuition of frequency-domain interference from Refs. [22, 23], we propose a concrete sample-efficient delay measurement approach, Algorithm 1, that provably uses as few photons as possible. In particular, letting $T$ be the upper limit of $\Delta t$, $t_{c}$ be the coherence time of the photon without lensing (we consider $t_{c}$ to be an inherent parameter of the photons for now; we will explain how $t_{c}$ is defined in a realistic observation scenario later in this paper), and assuming $\Delta t\gg t_{c}$, our method consumes only $O(\log(T/t_{c}))$ photons to measure $\Delta t$ with precision $t_{c}$, while traditional proposals require $O(T/t_{c})$ photons. Note that we work in the photon-starved regime, hence we expect to receive at most one photon per wave packet, and the wave function of each photon is the superposition of two wave packets separated by $\Delta t$ due to the microlensing effect. We also provide a rigorous proof that $\Omega(\log(T/t_{c}))$ is the information-theoretic lower bound, hence our method is optimal. One proof is based on modeling the gravitational lensing system as a communication channel and computing its channel capacity. We also exploit a surprising connection between the delay-finding problem and a well-studied problem in quantum computing, the dihedral hidden subgroup problem. We show that the dihedral hidden subgroup problem can be reduced to our problem, giving an alternative optimality proof in terms of both sample complexity and computational complexity.
@@
The key ingredient for the exponential improvement in our scheme is that our algorithm uses quantum information processing technologies (including single-photon spectrometers, and, depending on the specific implementation of our scheme, quantum memory and digital quantum computation) to perform single-photon frequency-basis measurements. This allows us to sample from a certain distribution determined by the value of $\Delta t$. By feeding these samples into a data-processing procedure inspired by the sample-efficient algorithm for the dihedral hidden subgroup problem [33], we can estimate $\Delta t$ in the style of maximum-likelihood estimation.
@@
Implementing our approach involves measuring the frequency of every photon, which requires a broadband high-resolution spectrometer with single-photon sensitivity. The difficulty of the implementation strongly depends on $T$, the upper limit of $\Delta t$ (which is determined by the lensing object) of our interest, because the required frequency resolution is $\sim 1/\Delta t$. Single-photon spectrometers based on dual-combs feature up to $\sim 100\,\mathrm{MHz}$ resolution [34, 35, 36, 37] with $\sim 10\,\mathrm{GHz}$ bandwidth. State-of-the-art spectrometers based on a time lens [38, 39, 38] even achieve $20\,\mathrm{kHz}$ resolution, but their bandwidth is limited to MHz-level. These results in principle enable measurements of $10\,\mathrm{ns}$ (dual-comb) or up-to-$0.1\,\mathrm{ms}$ (time lens) time delay, corresponding to lensing objects as heavy as brown dwarfs (dual-comb) or primordial black holes of multiple solar masses (time lens), as is explained later in Section II.1. However, one bottleneck for such an observation is that the photon sources we consider are thermal sources emitting broadband signals. The tiny bandwidth of existing single-photon spectrometers may require using prohibitively many of them in parallel. Therefore, long-$\Delta t$ measurements may only be achievable through next-generation single-photon spectrometers; however, for shorter $\Delta t$, which are also of significant interest to astronomy, resolving the frequency of single photons is much less challenging and can potentially be realized by combining existing technologies.
@@
We also propose another version of the delay-finding approach (Algorithm 2) which relies on storing and processing the photon wave function in the time domain and uses a different data processing procedure. We propose to first perform non-demolition frequency measurement on the received photon to localize it to a frequency range with width denoted by $1/t^{\prime}_{c}$ (where $t^{\prime}_{c}$ is the effective coherence time satisfying $t^{\prime}_{c}\geq t_{c}$). Then, using a quantum information discarding process, one can store the photon in a digital quantum computer in an undersampling manner in the time domain. More specifically, we propose to use a quantum memory that can only distinguish $O(T/t^{\prime}_{c})$ modes, which is far fewer than $\Theta(T\omega_{0})$ modes for sampling at the Nyquist rate (where $\omega_{0}$ is the carrier frequency of the photon). With this approach, one can employ the quantum Fourier transform (QFT) to produce an aliased frequency as the output, which is fed into Algorithm 2 to find $\Delta t$. The connection between delay finding and the dihedral hidden subgroup problem is established through this time-domain version. We show that this version can potentially be implemented by a linear optics system. Going further, digital quantum computing in principle enables compressed storage of the photonic modes with binary encoding of arrival time, which gives an exponential reduction in the resources required.
@@
Our photon-efficient method enables the estimation of microlensing time delays in the optical and infrared (IR) bands. The measurable delay range spans from $10^{-10}\,\mathrm{s}$ to $10^{-3}\,\mathrm{s}$ depending on the capabilities of the frequency-resolving device. This, in principle, supports observation of many more sources than in radio/gamma wavelengths, and the photons can be lensed by the majority of interesting microlensing systems. However, optical/IR waves oscillate extremely fast, giving a rather stringent requirement on the variance of $\Delta t$ between different photons due to the finite-source effect. This means it is much more difficult to have ingredient (ii). Nevertheless, as part of our solution, we give a concrete use case of our measurement scheme in the optical/IR band that satisfies ingredient (ii), and present a comprehensive analysis of its scientific value and feasibility. Specifically, we consider flares of M-class red dwarfs (M dwarfs). M dwarfs are relatively tiny and cold stars, and a flare is an event in which a small region of the dwarf becomes almost as bright as the whole dwarf in certain passbands. Our analysis shows that, for a significant fraction of flares in M dwarfs, the size of the light-emitting area may be small enough such that the uncertainty in $\Delta t$, denoted by $\delta_{\Delta t}$, is less than one period of the carrier frequency $\sim 10^{-15}$s. Moreover, our scheme not only enables the study of the lensing object, but also yields constraints on the actual spatial size of the flares in M dwarfs, which is currently poorly understood. Indeed, directly resolving flare kernels on even nearby M dwarfs likely requires $\geq 10\,\mathrm{km}$ optical baselines, which may only be achievable with quantum-assisted optical interferometry [4]. We also perform a comprehensive analysis of the number of photons we can receive in realistic settings to observe microlensed flares on M dwarfs, taking into account the duration, size, and temperature of the flare, as well as astronomical dust extinction and telescope collecting area. Our result shows that near-term ground-based optical telescopes can achieve sufficiently high signal-to-noise ratio in such observations. Moreover, we improve our algorithms such that photons from temporally and spatially separated flares can be analyzed collectively to infer the average lensing time delay, allowing for potential implementation using existing optical telescopes.
@@
Finally, we discuss the application of our methods to a different task: the calibration of time delays in telescope arrays. Light traveling from a source along different paths picks up a relative delay when observed at different sites. Learning these time delays is important to enable interference of photons arriving at different telescopes. In order to learn the time delays, we map the distributed problem to a lensing-like scenario, where a single detector observes the photons. In particular, we show how to use entanglement to transfer the information across the array to a single site, where we can apply our algorithm. The same compression and storage of photonic information in memory as used in prior work on telescope arrays [5, 6] is applicable here, such that our proposal is compatible with that scheme. The benefits over classical techniques are a replacement of long delay lines with memories that keep track of timing, allowing for longer baselines and thus larger resolution, and improved sample efficiency, as provided by our algorithm.
@@
The remainder of the paper is structured as follows. We provide preliminary information in Section II: we give a technical introduction to gravitational lensing and the significance of measuring its time delay (LABEL:{sec:lens_intro}), explain the setup of the delay-finding problem with mathematical and physical rigor (Section II.2), and review a traditional time delay measurement approach using a large number of photons (Section II.3). In Section III, we describe our sample-efficient delay-finding algorithm: we introduce the frequency-domain interference framework via a classical electromagnetism derivation (Section III.1), reproduce the same derivation in the quantum picture for photonic wave functions (Section III.2), use this picture to propose our delay-finding algorithm and analyze its sample complexity (Section III.3), explain the consequence of varied $\Delta t$ values (the finite-source effect) and analyze the how our algorithm performs under this effect (Section III.4), analyze the effect of noise photons (Section III.5) and the lensing magnification (Section III.6), and finally discuss the realistic scenario of broadband input photons which induces a short coherence time (Section III.7). In Section IV, we prove the information-theoretic lower bound for the sample complexity, matching the actual sample complexity of our algorithm. In Section V, we propose the undersampling algorithm using the quantum Fourier transform on a digital quantum computer, and discuss its connection to the dihedral hidden subgroup problem (Section V.3), giving another proof of the optimal sample complexity as well as computational complexity. In Section VI, we discuss possible experimental realizations of our algorithms. In Section VII, we present the astronomical observation plan: we carry out a case study for the example setup for microlensed M dwarf flares with an analysis of its feasibility and scientific value (Section VII.1 and a narrow-band version in Appendix B); we then introduce a modified version of our algorithm to combine photons from different flares and present numerical simulation results (Section VII.2). In Section VIII, we perform a thorough analysis for the robustness of our approach to noises due to the medium between the source and the telescope: we investigate the effect of dust extinction (Section VIII.1 and a detailed proof in Appendix A), astronomical scintillation (Section VIII.2), and atmospheric fluctuation (Section VIII.3). In Section IX, we discuss the application of our delay-finding protocol to calibrating quantum telescope arrays. Finally, we summarize our work and discuss open problems in Section X.